3. A.

The disposable income of Mehta family increases from Rs 5000 to Rs 15,000. As a result,

the family‘s demand for milk and milk goods has increased from 30 liters to 60 liters per

month. Calculate the income elasticity of demand. (5 Marks)

3. B.

A drop in the price of lemons from Rs 100 per kg to Rs 60 Per Kg increases the quantity

demanded from 1.75 to 7 kg per week. Calculate the price elasticity of demand. (5 Marks)


1
Expert's answer
2021-06-29T19:33:02-0400

3.a)  income elasticity of demand= "\%" "change in quantity demanded\over change in icome"

percentage change of quantity demanded"=new value-old value\over old value"

percentage change of quantity demanded ="60-30\over30" =1

percentage change in income ="15000-5000\over5000" =2

Therefore price elasticity of demand ="1\over 2" =0.5

=0.5

3.b) price elasticity of demand = "percentage change in quantity demanded\over percentage change in price"

percentage change in quantity demanded ="7-1.75\over 1.75" =3

percentage change in price="60-100\over100" =0.4

Thus, price elasticity of demand ="0.4\over 3"

=0.1